James is four years younger than Austin. If three times James' age is increased by the square of Austin's age, the result is 28. Find the ages of James and Austin. (Algebraic solutions only)
Answer by scott8148(6628)

"James is four years younger than Austin" ___ j = a - 4

"If three times James' age is increased by the square of Austin's age, the result is 28"
___ 3j + a^2 = 28

substituting ___ 3(a - 4) + a^2 = 28 ___ a^2 + 3a - 40 = 0 ___ (a + 8)(a - 5) = 0

solve for a, then substitute back to find j